HUBBLE SPACE TELESCOPE DAILY REPORT #1303 PERIOD COVERED: 0000Z Daily Status Report as of 011/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: The science recovery SMS successfully started at 010/1200Z. No science observations were taken during the first 12 hours of this day. 1.1 Completed Three Sets of WF/PC-2 5562 (Cycle 4 Calibration: Internal Monitor 3: Darks) The first set of observations fell during the software safemode period, and no observations were taken. For the next two sets of observations, the WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5397 (Determination of the Extra galactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-3351. The images will be used in measuring the periods of Cepheid variable stars within the galaxies. The observations were completed as planned, and no anomalies were noted. 1.3 Completed FGS 5336 (Linking the HIPPARCOS to the Extra galactic Reference Frame: NEWHIP-B) The Fine Guidance Sensor #3 was used to make astrometric observations of objects NEWEGOB-20, 320-REF, 321-REF, 322-REF, NEWHIP-20. The observations were executed as scheduled. No anomalies were reported. 1.5 Completed FOS/RD 5680 (Supernova Remnants and Nucleosynthesis) The FOS (Red Side) was used to obtain spectra of supernova remnant 0504-69 in the LMC. There were four reported recentering events which lasted for a total of 19 seconds. 1.6 Completed WF/PC-2 5091 (GTO High Latitude Parallel Program - Cycle /4) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations greater than +/-30 degrees. One data group in one observation was lost during a tape recorder track change. This was an expected loss. Also, there was one reported recentering event which lasted for a total of four seconds. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Notes: The DF-224 Solar Array Profile Limit Test was changed from 14 to 10 degrees. Data inversions while utilizing STGT continue. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-